OpenStudy (elisaneedshelp):

What is the slope of the line given by the equation y + 2 = −3(x − 7)? A. m = −3 B. m = 3 C. m = 2 D. m = −7

OpenStudy (anonymous):

B

jigglypuff314 (jigglypuff314):

B is incorrect point slope form: y - y1 = m(x - x1) when m = slope and given the point (x1, y1) in y + 2 = −3(x − 7) ^^ is the slope

OpenStudy (anonymous):

-3

OpenStudy (elisaneedshelp):

im guessing -3 :P